Find f '(1).
f(x) = (ln(x)) / x^9
1
Expert's answer
2011-02-28T05:16:44-0500
f'(x) = 1/x * 1/x9 - 9 ln(x)/x10 = 1/x10 (1- 9ln(x))
f'(x) = 1/110 (1 - 9 ln(1)) = 1 (1-0) = 1